OpenStudy (anonymous):

what is the simplified form of y^2+6y+8/y^2+7y+12 ?

OpenStudy (anonymous):

\[\frac{y^2+6y+8}{y^2+7y+12}=\frac{(y+4)(y+2)}{(y+4)(y+3)}=\frac{y+2}{y+3}\]

OpenStudy (anonymous):

thank you. :)